OpenStudy (anonymous):

A basketball is dropped from an original height of 9 feet and on each bounce, bounces up 2/3 the distance it fell. How far will it have traveled by the time it hits the ground the 4th time?

OpenStudy (anonymous):

It is a G.P....With a = 9 r = 2/3 Sn = a ( 1-r^n) / (1-r)

OpenStudy (anonymous):

:) enjoy

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

what would n be? or am I solving for n?

OpenStudy (anonymous):

S_4 = 2.407?